# Testing strong-field QED with the avalanche precursor

### v1

@@ IV Parametric study
Figure 8: Final total energy of photons and electrons as a function of the peak total power (or the total field strength, shown on the upper horizontal axis) for the case of LP $f/2$. The data is normalized by the total number of seed electrons in the simulation box $N_{\mathrm{seed}}\approx 4\times 10^{7}$.
@@ IV.1 Idealized seed electron target
For this study, we use an idealized target for injection of electrons in the focus. We assume that it consists only of electrons homogeneously distributed in a ball of radius $1.34$ $\mu$m. In our simulations, we build on the setup from Sec. III.2 and vary its parameters. To speed up the numerical computation, we start from the time $t=-0.5T$, assuming that the laser fields have already arrived at the interaction point (recall that the maximum electric field reaches the origin $\mathbf{r}=\mathbf{0}$ at $t=0$). The electrons are injected in the field at the start of the simulation. Finally, we set the particle density to $n=3.97\times 10^{18}$ cm-3. With this density, the number of seed electrons in the small focal volume $\lambda^{3}$ matches $N_{\mathrm{seed}}^{\mathrm{foc}}$ for the krypton target.
@@ IV.2 Total particle number and energy dependence on the laser parameters
Let us consider the dependence of the total number of positrons and high-energy ($>1$ MeV) photons (for brevity, the signal) produced during the interaction. In Fig. 7 we show its dependence on the total power $P_{*}$ delivered by two laser pulses. Note that while we reference the electron density to the krypton gas target, these results can be rescaled (at least within the range of densities for which collective effects are negligible), therefore, it is convenient to normalize the signal by $N_{\mathrm{seed}}^{\mathrm{foc}}$.
@@
The transitions between the interaction regimes can be clearly distinguished in the field strength dependence of the total energy contained in photons and electrons after the interaction. We plot this quantity in Fig. 8 for the case of LP and $f/2$ focusing for (i) all particles in the simulation box and (ii) particles that propagate along the electric field polarization axis in the angle opening $\pi/2-\Delta\theta<\theta<\pi/2+\Delta\theta$, $\Delta\theta=0.15\pi$ (as with the orange curves in the spectral plots in Fig. 5). The total energy is calculated as the average energy (extracted from the spectral angular-resolved data) multiplied by the corresponding number of particles. We choose to normalize the curves by the total number of seed electrons available in the system at the start of the simulation.
@@
To summarize, an analysis of the particle numbers, energies and spectra as a function of the laser power and intensity can be used to identify the regime of interaction. Furthermore, a study of the distributions for particles moving in the direction of the electric field polarization can be used to deduce most of the characteristics of the interaction dynamics.
@@ V Realistic laser pulse alignment and desynchronization
$30^{\circ}$ $\Delta y=1.5$ $\mu$m - 0.48 0.57 2.8
$30^{\circ}$ $\Delta z=1.5$ $\mu$m - 0.27 0.53 2.8
$30^{\circ}$ $\Delta z=1.5$ $\mu$m 10 fs 0.21 0.45 2.0
@@
Recall that in the ideal case, the laser pulses exactly counterpropagate with respect to each other along the $x$-axis, are polarized linearly in the $y$-direction and focused to $f/2$ (as in Section III). Then, the laser propagation axes are tilted in the $(xy)$-plane, so that the relative angle between them amounts to 30 (in simulations, we tilt each beam by 15, as this angle allows avoiding back-propagation at $f/2$ focusing). To account for possible mismatch of the focal centers (for brevity, “misfocus”), we shift their position in the $(yz)$-plane at $x=0$. We tested two cases, when the focal centers are shifted in the $y$- or $z$-direction by 1.5 $\mu$m. Finally, we assume that one of the pulses can be delayed by 10 fs (“mistiming”).
In this set of simulations, to treat the electron injection realistically, we consider a Kr26+ gas target. It is shaped as a rectangular block sized $d_{x}=0.75w_{0}=1.34$ $\mu$m along the $x$-axis, and $d_{y}=d_{z}=2d_{x}=2.68$ $\mu$m in the transverse directions. [^12] This change from the ball-shaped considered in Sections III and IV is motivated by the convenience for studying the effects of misfocusing. As in Section III, at the start of each simulation, the laser pulses are injected from the simulation box sides and propagate to the center, ionize the target and trigger a cascade. The simulation parameters are also the same as described in Section III.2.
@@ Acknowledgements
We express our gratitude to J. D. Zuegel for discussions and support of this project.
@@
All of the simulation results we obtained on the HPC system “3Lab Computing” hosted at École polytechnique and administrated by the Laboratoire Leprince-Ringuet, Laboratoire des Solides Irradiés and Laboratoire pour l’Utilisation des Lasers Intenses. S.S.B. was supported by U.S. Department of Energy Office of Science Office of High Energy Physics under Contract No. DE-AC02-05CH11231. A.D.P. is partially supported by the U.S. National Science Foundation Mid-scale Research Infrastructure Program under Award No. PHY-2329970. The work of G.G. was partially supported by UKRI under grants ST/W000903/1 and EP/Y035038/1.
@@ Appendix A The probability rates of the nonlinear Compton emission and Breit-Wheeler pair production
[…]
@@
[…]
@@
[…]
@@
$W_{\mathrm{nBW}}(\chi_{\gamma},\gamma_{\gamma})\approx\frac{\alpha}{\tau_{C}\gamma_{\gamma}}\times\left\{\begin{array}[]{ll}0.23\,\chi_{\gamma}e^{-\frac{8}{3\chi_{\gamma}}},&\chi_{\gamma}\ll 1,\\ 0.38\,\chi_{\gamma}^{2/3},&\chi_{\gamma}\gg 1.\end{array}\right.$ (16)

### v2

@@ IV Parametric study
Figure 10: Final total energy of photons and electrons as a function of the peak total power (or the total field strength, shown on the upper horizontal axis) for the case of LP $f/2$. The plotted energy is per seed electron ($N_{\mathrm{seed}}\approx 4\times 10^{7}$).
@@ IV.1 Idealized seed electron target
[…]
@@ IV.2 Total particle number and energy dependence on the laser parameters
Let us consider the dependence of the total number of positrons and high-energy (>1>1 MeV) photons (for brevity, the signal) produced during the interaction. In Fig. 9 we show its dependence on the total power $P_{*}$ delivered by two laser pulses. Note that while we reference the electron density to the krypton gas target, these results can be rescaled (at least within the range of densities for which collective effects are negligible), therefore, it is convenient to normalize the signal by $N_{\mathrm{seed}}^{\mathrm{foc}}$.
@@
The transitions between the interaction regimes can be clearly distinguished in the field strength dependence of the total energy carried by photons and electrons after the interaction. We plot this quantity in Fig. 10 for the case of LP and $f/2$ focusing for (i) all particles in the simulation box and (ii) particles that propagate along the electric field polarization axis in the angle opening $\pi/2-\Delta\theta<\theta<\pi/2+\Delta\theta$, $\Delta\theta=0.15\pi$ (as with the orange curves in the spectral plots in Fig. 7). [^12] We use the following procedure to calculate the total energy carried by particles moving in a given interval of angles $\theta_{1}<\theta<\theta_{2}$. We extract the energy-angular distribution $f(\varepsilon,\theta)$ from simulations at a time point after the interaction. Given this distribution is normalized such that the integration gives the number of particles contained in the chosen part of the phase space, $N=\int_{\theta_{1}}^{\theta_{2}}d\theta\,\sin\theta\int d\varepsilon\,f(\varepsilon,\theta)$, the energy carried by these particles can be calculated as $\varepsilon_{\mathrm{tot}}=\int_{\theta_{1}}^{\theta_{2}}d\theta\,\sin\theta\int d\varepsilon\,\varepsilon f(\varepsilon,\theta)$. In Fig. 10, we plot $\varepsilon_{\mathrm{tot}}/N_{\mathrm{seed}}$, namely, energy per seed electron.
@@
To summarize, an analysis of the particle numbers, energies and spectra as a function of the laser power and intensity can be used to identify the regime of interaction. Furthermore, a study of the distributions for particles moving in the direction of the electric field polarization can be used to analyse the interaction at the focal center. For example, the spectral cut-off location in this distribution can be employed to discriminate between the classical and quantum RD regimes. The dependence of the signal (from particles collected in the transverse direction) on the field intensity can be used to identify the transition to an avalanche. Finally, the shape of the spectra can also be used to deduce the cascade parameters, such as energy and $\chi$ of particles, by fitting analytical models.Nerush, Bashmakov, and Kostyukov (2011)
@@ V Realistic laser pulse alignment and desynchronization
[…]
@@
[…]
@@ Acknowledgements
We express our gratitude to J. D. Zuegel for discussions and support of this project. We are also grateful to Jake Bromage for providing useful information on the envisaged contrast at NSF OPAL.
@@
All of the simulation results we obtained on the HPC system “3Lab Computing” hosted at École polytechnique and administrated by the Laboratoire Leprince-Ringuet, Laboratoire des Solides Irradiés and Laboratoire pour l’Utilisation des Lasers Intenses. S.S.B. was supported by U.S. Department of Energy Office of Science Office of High Energy Physics under Contract No. DE-AC02-05CH11231. A.D.P. is partially supported by the U.S. National Science Foundation Mid-scale Research Infrastructure Program under Award No. PHY-2329970. The work of G.G. was partially supported by UKRI under grants ST/W000903/1 and EP/Y035038/1. S. Meuren is supported by the Centre National de la Recherche Scientifique (CNRS) and the Agence Nationale de la Recherche (ANR) under the Chaire de Professeur Junior (CPJ) program. Financial support by the ANR (g4QED project, Grant No. ANR-23-CE30-0011) is acknowledged.
@@ Appendix A The probability rates of the nonlinear Compton emission and Breit-Wheeler pair production
$\frac{dW_{\mathrm{nCE}}(\gamma_{e},\chi_{e},\gamma_{\gamma})}{d\xi}=\frac{1}{\sqrt{3}\pi}\dfrac{\alpha}{\tau_{C}\gamma_{e}}\left[\left(1-\xi+\frac{1}{1-\xi}\right)K_{2/3}(\mu)-\int_{\mu}^{\infty}dsK_{1/3}(s)\right],$ (8)
@@
$\frac{dW_{\mathrm{nBW}}(\gamma_{\gamma},\chi_{\gamma},\gamma_{e})}{d\zeta}=\frac{1}{\sqrt{3}\pi}\dfrac{\alpha}{\tau_{C}\gamma_{\gamma}}\left[\left(\frac{\zeta}{1-\zeta}+\frac{1-\zeta}{\zeta}\right)K_{2/3}(\mu^{\prime})-\int_{\mu^{\prime}}^{\infty}dsK_{1/3}(s)\right],$ (9)
@@
$W_{\mathrm{nCE}}(\gamma_{e},\chi_{e})=\int_{0}^{1}\frac{dW_{\mathrm{nCE}}(\gamma_{e},\chi_{e},\gamma_{\gamma})}{d\xi}d\xi,$ (10)
$W_{\mathrm{nBW}}(\gamma_{\gamma},\chi_{\gamma})=\int_{0}^{1}\frac{dW_{\mathrm{nBW}}(\gamma_{\gamma},\chi_{\gamma},\gamma_{e})}{d\zeta}d\zeta$ (11)
@@
$W_{\mathrm{nBW}}(\gamma_{\gamma},\chi_{\gamma})\approx\frac{\alpha}{\tau_{C}\gamma_{\gamma}}\times\left\{\begin{array}[]{ll}0.23\,\chi_{\gamma}e^{-\frac{8}{3\chi_{\gamma}}},&\chi_{\gamma}\ll 1,\\ 0.38\,\chi_{\gamma}^{2/3},&\chi_{\gamma}\gg 1.\end{array}\right$. (17)
@@
Appendix B Approximations in use for modelling elementary SFQED processes and avalanche-type cascades
The kinetic approach underlying PIC-MC simulations is based on the interplay of several approximations.Gonoskov et al. (2015) As the laboratory-frame fields that we consider are below the Sauter-Schwinger limit, $E\ll E_{S}$, we can describe particle dynamics semi-classically. At the same time, the studied fields are strong in the sense that $a_{0}\gg 1$ (for example, $a_{0}\gtrsim 5\times 10^{2}$ per beam at NSF OPAL). We consider a standing wave, and the electric field in the magnetic node acts twofold: (i) it accelerates charged particles and provides high $\gamma_{e}$ at a time $1/\omega$, and (ii) it induces the quantum processes. The most relevant ones are the photon emission and pair creation, and their formation scale is typically of the order $\sim O(\lambda/a_{0})\ll\lambda$. This enables us to use the locally constant field approximation (LCFA), see the reviews Di Piazza et al. (2012); Gonoskov et al. (2022); Fedotov et al. (2023); Popruzhenko and Fedotov (2023) for a detailed discussion, and also Refs.Ritus (1985); Baier and Katkov (1968) Indeed, in the kinetic picture, a particle propagates along a classical trajectory until it undergoes a quantum process randomly.
Let us discuss the set of used approximations in more detail. First, recall that a QED cascade is a high-multiplicity process. As the first major simplification, it is customary to consider it as a chain of first-order processes, namely, the nonlinear Compton emission and Breit-Wheeler pair creation. Such factorization is well-justified for processes in high fields with $a_{0}\gg 1$ because the formation length, which scales as $l_{f}\sim\lambda/a_{0}$ is much smaller than a laser period and then than the pulse length $c\tau$. Thus, a factorized process made of $n$ elementary processes scales as $(c\tau/l_{f})^{n}\sim(c\tau a_{0}/\lambda)^{n}$, whereas a nonfactorized process is linear in $c\tau$ and thus roughly scales as $c\tau a_{0}/\lambda$.Baier and Katkov (1968) Hence, at high $a_{0}$, the former channel dominates, which motivates the chain approximation.
Second, let us briefly discuss the LCFA. This is a common approximation for the emission and pair creation rates when modelling cascades. Within it, the rates are defined by the local values of the background field at the position of the incoming particle. Due to the relativistic effects, the field in the particle rest frame is approximated by a constant crossed field (CCF), $E=B$, $\mathbf{E}\perp\mathbf{B}$, which enters the probabilities $W_{\mathrm{nCE,nBW}}$ only through the dependence on $\chi_{e,\gamma}$, see Eqs. (8), (9).
A rigorous formulation of the LCFA applicability conditions is a challenging problem in SFQED, and we are not aware of a derivation valid in a general background field. For photon emission and pair creation, the LCFA rates typically arise in the limit of high $a_{0}$ and $\gamma_{e}$, which was shown for a magnetic field,Baier and Katkov (1968) a plane wave,Nikishov and Ritus (1964); Di Piazza et al. (2018, 2019); Ilderton, King, and Seipt (2019); Heinzl, King, and MacLeod (2020); King (2020) a short laser pulse,Seipt and Kämpfer (2011); Ilderton (2019); Titov, Otto, and Kämpfer (2020); Kämpfer and Titov (2021) focused beam,Di Piazza (2021); Nielsen, Holtzapple, and King (2022); Adamo and Ilderton (2025) crystal field, Wistisen and Di Piazza (2018, 2019) or in a general time-dependent electric field.Gelfer et al. (2022); Mironov, Gelfer, and Fedotov (2023) As we consider the interaction at the magnetic node of a standing wave, the latter field model is relevant for our case. Following Ref.Gelfer et al. (2022), for photon emission, the LCFA applicability requires:
  1. 1.
    Locality — $a_{0}\gg 1$ and $a_{0}^{3}\gg\chi_{e}\chi_{e}^{\prime}/\chi_{\gamma}$.
  2. 2.
    The field is close to crossed in the incoming and outgoing particles’ frame — $\gamma^{3}\gg\chi_{e}\chi_{e}^{\prime}/\chi_{\gamma}$.
Notably, the LCFA is a general approximation for SFQED effects not specific to photon emission and pair creation, and could be used, e.g. to treat Sauter-Schwinger pair creation in standing waves,Aleksandrov, Plunien, and Shabaev (2019); Kohlfürst et al. (2022) however, the applicability conditions can be different.
Generally speaking, the corrections to the LCFA can be relevant, for example, at $a_{0}\approx 1-10$, and $\gamma_{e,\gamma}\sim 10^{2}$. Such corrections can be calculated numerically, e.g. for a plane wave background or a single laser pulse.Di Piazza et al. (2019); Blackburn, MacLeod, and King (2021); Blackburn and King (2022); Montefiori and Tamburini (2023) At higher $a_{0}$ and $\gamma_{e,\gamma}$, the LCFA can be considered universally robust for various background field models.
When modelling the onset stage of avalanche-type cascades, it is customary to consider the seed electrons initially at rest injected in the magnetic node of a standing wave. One could argue that for such initial conditions, the LCFA rates may require corrections when describing photon emission (see e.g. Refs. Raicher, Eliezer, and Zigler (2015); King and Hu (2016) for details). At $a_{0}\gg 1$, this applies if $\gamma_{e}$ for the emitting electron is relatively low (see the second condition of the applicability of the LCFA). Let us compare the time scales of emission and electron acceleration for the parameters of interest here. During acceleration by the field, the electron gains energy $\sim mc^{2}$ at a time scale $1/\omega a_{*}\lesssim t_{\mathrm{rel}}\ll 1/\omega$. The characteristic time, at which it emits a photon, can be estimated from the LCFA rate: $t_{\mathrm{em}}\sim W_{\mathrm{nCE}}^{-1}(\gamma_{e}(t_{\mathrm{em}}),\chi_{e}(t_{\mathrm{em}}))$ [see also Eq. (4) and Ref.Mercuri-Baron et al. (2025)]. Assuming that the emission takes place at $\chi_{e}\lesssim 1$ (at $\chi_{e}\gg 1$ the LCFA is typically valid), the low-$\chi_{e}$ asymptotic of Eq. (14) gives $t_{\mathrm{em}}\sim 1/\sqrt{\alpha a_{*}}\omega$. Hence, $t_{\mathrm{em}}/t_{\mathrm{rel}}\sim\sqrt{a_{*}/\alpha}\gg 1$ for $a_{*}\sim 10^{3}$. Therefore, the electron gains high $\gamma_{e}$ during the phase of semi-classic motion before the instance of emission, which, in turn, can be safely treated within the LCFA. We note that the robustness of the LCFA should be verified with a different approach if $\omega t_{\mathrm{em}}\gtrsim 1$. However, this case does not fall within the scope of the current work.
It should also be noted that the LCFA is known to fail in the infrared part of the radiation spectrum.Di Piazza et al. (2018, 2019); Ilderton, King, and Seipt (2019) However, soft photons do not contribute to the formation of cascades, as such photons do not create pairs and do not cause significant recoil. Hence, the region of the parameter space where LCFA fails is irrelevant for modelling avalanche-type cascades.
Finally, when higher-order processes in a cascade (such as tridents) are factorized, the spin states of intermediate particles should be treated as well. The common approach is taking spin-averaged probability rates for the first-order processes [given in Eqs. (8) and (9)]. A more rigorous approach suggests using spin-resolved rates.King, Elkina, and Ruhl (2013); Seipt et al. (2021) It is possible to include such rates explicitly in the kinetic descriptionSeipt and Thomas (2023) and simulation codes.Zhao et al. (2023); Qian et al. (2023) The main effect is the cascade growth rate dampening. However, simulations show that for background fields of optical frequency, calculations with spin-averaged rates provide a reliable result for the avalanche growth rate (we also assume that seed electron states are averaged over their initial states).King, Elkina, and Ruhl (2013); Seipt et al. (2021) In our simulations with SMILEI,Derouillat et al. (2018) we rely on spin-averaged rates for the quantum processes. Still, it can be interesting to include spin dynamics in future works to identify its effect on final particle distributions in the parameter range of future experiments.
A more rigorous derivation of all the listed approximations to modelling avalanches within the kinetic approach would be undoubtedly of great interest and should be done in future, for example, with the techniques developed in Ref.Fauth, Berges, and Di Piazza (2021)
Appendix C Dependence of the final particle spectra on the angular spread
Figure 12: The final particle spectra. The black, orange, and darker blue curves, as well as the vertical lines, are taken from Fig. 7. The spectra of particles propagating in the transverse direction $\pi/2-\Delta\theta<\theta<\pi/2+\Delta\theta$ (denoted $\perp$” in the legend) are shown in warm colors for different values of the opening angle $\Delta\theta$. Similarly, curves in cold colors correspond to particles propagating longitudinally $0<\theta<\Delta\theta$ (denoted $\parallel$” in the legend).
In Section III.5, we introduced the angular spread $\Delta\theta$ to separate the dynamics of particles propagating in the transverse direction $\pi/2-\Delta\theta<\theta<\pi/2+\Delta\theta$ and in the longitudinal direction $0<\theta<\Delta\theta$. As we discussed, particles propagating transversally are associated with the avalanche-type cascade, and one can study, for example, their spectra selectively, as shown in Fig. 7. In Section III.5 and later in Fig. 10, we set $\Delta\theta=0.15\pi$, as this value is convenient for illustration. For example, this can be seen from the green curve in Fig. 8e showing the distribution of positrons in $\theta$. We associate the pronounced hump centered at $\theta=\pi/2$ to particles propagating directly from the interaction point, where the cascade is developing. The smaller peaks near $\theta\approx 0$ and $\theta\approx\pi$ correspond to particles accelerated by each of the laser pulses, as they propagate after the impact. We find the spread $\Delta\theta=0.15\pi$ representative to discriminate between the longitudinal and transverse motion.
Here, we show the dependence of the final particle spectra on $\Delta\theta$, which is presented in Fig. 12. As one can see, the overall shape of the spectra is stable at varied $\Delta\theta$. In particular, the position of the high-energy cut-off for electrons and positrons propagating transversely matches for different $\Delta\theta$. At the largest spread, $\Delta\theta=0.25\pi$, the longitudinal distribution (the purple line) becomes noticeably wider in the lower energy. This motivates considering narrower angular spreads for the extraction of cleaner signals. Overall, our discussion and conclusions presented above are not sensitive to variation of $\Delta\theta$ within reasonable limits (namely, not too small so that there is enough statistics for a measurement, and not too large, so that the transverse and longitudinal motion are separated).